# Scalable multiphoton generation from cavity-synchronized single-photon sources

### v1

@@ VI Conclusions and Outlook
The realization of scalable and deterministic sources of multi-photon states opens the door to large-scale quantum information processing with photons such as boson sampling Brod et al. (2019), but also quantum networking Llewellyn et al. (2020), simulation Sparrow et al. (2018), and computation Takeda and Furusawa (2019). Moreover, the setup and ideas introduced in this work can be further extended to generate correlated multi-photon states Meyer-Scott et al. (2019); Pfaff et al. (2017) with engineered entanglement patterns. Many-body methods such as Matrix-product-states Huang et al. (2019); Orús (2014) could be exploited to study the multi-photon correlations of the propagating fields, and recent multi-photon probing methods Ramos and García-Ripoll (2017); Sánchez-Muñoz et al. (2018); Dassonneville et al. (2020b); Lescanne et al. (2020); Jeannic et al. (2020) could be use to characterize them in the lab.
@@ A.1 Coherent displacement of driven cavity mode
$\dot{a}={}$ $-\left[\kappa/2-i\Delta\right]a-i\sum_{j=1}^{N}g_{j}\sigma_{j}^{-}-i\Omega-\sqrt{\kappa}a_{\rm in},$ (36)
$\dot{\sigma}_{j}^{-}={}$ $-\left[(\gamma_{j}+\gamma^{j}_{\rm loss})/2-i\delta_{j}\right]\sigma_{j}^{-}+ig_{j}a\sigma_{j}^{z}$ (37)
$+\sqrt{\gamma_{j}}\sigma_{j}^{z}b_{\rm in}^{j}+\sqrt{\gamma^{j}_{\rm loss}}\sigma_{j}^{z}c_{\rm in}^{j}-i\sqrt{2\gamma_{\phi}^{j}}\chi(t)\sigma_{j}^{-},$
$\dot{\sigma}_{j}^{z}={}$ $-(\gamma_{j}+\gamma^{j}_{\rm loss})(1+\sigma_{j}^{z})+2ig_{j}(a^{{\dagger}}\sigma_{j}^{-}-\sigma_{j}^{+}a)$ (38)
$-2\sqrt{\gamma_{j}}\left[\sigma_{j}^{+}b_{\rm in}^{j}+{\rm h.c.}\right]-2\sqrt{\gamma^{j}_{\rm loss}}\left[\sigma_{j}^{+}c_{\rm in}^{j}+{\rm h.c.}\right].$
@@
$\delta\dot{a}={}$ $-\left[\kappa/2-i\Delta\right]\delta a-i\sum_{j=1}^{N}g_{j}\sigma_{j}^{-}-\sqrt{\kappa}a_{\rm in}(t),$ (39)
@@
$\dot{\sigma}_{j}^{-}={}$ $-\left[(\gamma_{j}+\gamma^{j}_{\rm loss})/2-i\delta_{j}\right]\sigma_{j}^{-}+ig_{j}\delta a\sigma_{j}^{z}+ig_{j}|\alpha|\sigma_{j}^{z}$
$+\sqrt{\gamma_{j}}\sigma_{j}^{z}b_{\rm in}^{j}+\sqrt{\gamma^{j}_{\rm loss}}\sigma_{j}^{z}c_{\rm in}^{j}-i\sqrt{2\gamma_{\phi}^{j}}\Theta(t)\sigma_{j}^{-},$ (40)
$\dot{\sigma}_{j}^{z}={}$ $-(\gamma_{j}+\gamma_{\rm loss}^{j})(1+\sigma_{j}^{z})+2ig_{j}(\delta a^{{\dagger}}\sigma_{j}^{-}-\sigma_{j}^{+}\delta a)$
$-2ig_{j}|\alpha|(\sigma_{j}^{+}-\sigma_{j}^{-})-2\sqrt{\gamma_{j}}\left[\sigma_{j}^{+}b_{\rm in}^{j}(t)+{\rm h.c.}\right]$
@@
Here, $\alpha(t)$ is determined via the classical differential equation (8). Notice that, without loss of generality, we have absorbed the phase of $\alpha$ in the definition of all raising and lowering operators, so that only the magnitude $|\alpha|$ has a physical effect in the dynamics of the Langevin equations (39)-(41).
@@
$H^{\prime}(t)={}$ $-\Delta\delta a^{{\dagger}}\delta a-\frac{1}{2}\sum_{j=1}^{N}\delta_{j}\sigma_{j}^{z}+\sum_{j=1}^{N}g_{j}(\delta a^{{\dagger}}\sigma_{j}^{-}+\sigma_{j}^{+}\delta a)$
$+\sum_{j=1}^{N}g_{j}|\alpha|(\sigma_{j}^{+}+\sigma_{j}^{-}).$ (43)
Notice that the master equation (42) for the dynamics of $\delta a$ is the same as the master equation (6) for the total field $a$, except for the driving term in displaced Hamiltonian (43), which acts on the qubits rather than on the cavity, i.e.
$\Omega(t)(a+a^{\dagger})\rightarrow\sum_{j=1}^{N}g_{j}|\alpha|(\sigma_{j}^{+}+\sigma_{j}^{-}).$ (44)
This is the effective cavity-mediated driving on the qubits that allows for the synchronization of many SPSs in our scheme.
Another important advantage of the displaced master equation (42) is its efficiency for numerical simulations in the case of strong cavity driving $|\Omega(t)|\gg\kappa,|\Delta|$, where most of the cavity photons can be taken into account by the classical field $\alpha(t)$, and the fluctuation $\langle\delta a^{\dagger}\delta a\rangle$ is very weakly populated. This largely reduces the Hilbert space dimension needed to numerically simulate the dynamics of the driven cavity mode coupled to all qubits.
@@ A.2 Adiabatic elimination of cavity fluctuations
In the bad cavity limit $g_{j},|\delta_{j}|\ll\kappa$, the cavity reaches very quickly a coherent steady state $\alpha_{\rm ss}=-i\Omega/(\kappa/2-i\Delta)$, with the quantum fluctuations $\delta a$ close to the vacuum state $|0\rangle$. In this situation it is convenient to adiabatically eliminate the cavity fluctuation and obtain an effective dynamics for the degrees of freedom of the qubits only. To do so, we formally integrate Eq. (39) and apply the Markov approximation provided the qubits’ dynamics evolves slowly on the time-scale $\sim 1/\kappa$. As a result, we get
$\delta a(t)\approx-i\sum_{j}\frac{g_{j}}{(\kappa/2-i\Delta)}\sigma_{j}^{-}(t)-\frac{\sqrt{\kappa}}{(\kappa/2-i\Delta)}a_{\rm in}(t).$ (45)
Replacing the above expressions in the original quantum Langevin equations (37)-(38), we obtain the effective dynamics of the qubits, given by
@@
$-2\sqrt{\gamma^{j}_{\rm loss}}\left[\sigma_{j}^{+}c_{\rm in}^{j}+{\rm h.c.}\right],$ (47)
where the effective cavity-mediated decay $\gamma_{\rm cm}^{j}$, detuning $\delta_{\rm cm}^{j}$, hopping $J_{\rm cm}^{jl}$, and driving strength $\Omega_{\rm cm}^{j}$ are defined in Eqs. (10) and (14)-(16). Without loss of generality, we have reabsorbed the constant phase $ie^{i\Theta}$, with $\Theta={\rm Arctan}(2\Delta/\kappa)$ in all annihilation operators.
Finally, we can find a master equation which leads to the same effective Langenvin equations in Eqs. (46)-(47). This effective master equation after adiabatic elimination of the cavity mode is given in Eqs. (12)-(13).
@@ Appendix B Photon counting and calculation of photon generation probabilities
Photon detection and counting lies at the heart of Quantum Optics Gardiner and Zoller (2004); Plenio and Knight (1998) and the study of light-matter interactions Meystre (2006); Obukhov et al. (2012), and thus plenty of methods have been developed over time. In this Appendix we describe two methods for quantifying the efficiencies and photon statistics from the output of many SPSs. In Sec. B.1 we introduce an original photon counting method that we developed based on extending the master equation formalism. Then, in Sec. B.2, we discuss a more standard photon counting method using the quantum trajectory (QT) approach. This is useful when the Hilbert space of the system becomes very large, but at the expense of losing precision in the computation of the averages.
@@ B.2 Photon counting in quantum trajectories formalism
The most natural way to implement an ideal photon counting is within the formalism of quantum trajectories (QT) and continuous measurements Plenio and Knight (1998); Daley (2014); Dalibard et al. (1992); Molmer et al. (1993); Dum et al. (1992). Here, the physics of quantum jumps is explicitly simulated during the open system evolution and therefore it is very natural to count them and thereby infer the photon statistics.

### v2

@@ VI Conclusions and Outlook
Scalable and deterministic sources of multi-photon states will be a key resource for realizing large-scale quantum information processing such as quantum optical neural networks Steinbrecher et al. (2019) or fault-tolerant photonic quantum computation Takeda and Furusawa (2019). In the short-term, the implemenation of our scheme can already enable quantum advantage experiments with hundreds of microwave photons such as boson sampling Brod et al. (2019) or quantum metrology Motes et al. (2015); Ge et al. (2018). Moreover, the setup and ideas introduced in this work can be further extended to scale up the generation of correlated multi-photon states Meyer-Scott et al. (2019); Pfaff et al. (2017) with engineered entanglement patterns Besse et al. (2020). Many-body methods such as Matrix-product-states Huang et al. (2019); Orús (2014) could be exploited to study the multi-photon correlations of the propagating fields, and recent multi-photon probing methods Ramos and García-Ripoll (2017); Muñoz et al. (2018); Dassonneville et al. (2020b); Lescanne et al. (2020); Le Jeannic et al. (2021) could be use to characterize them in the lab.
@@ A.1 Coherent displacement of driven cavity mode
$\dot{a}={}$ $-\left[\kappa/2+i\omega_{c}\right]a-i\sum_{j=1}^{N}g_{j}(\sigma_{j}^{-}+\sigma_{j}^{+})-\sqrt{\kappa}a_{\rm in}$
$-2i\Omega(t)\cos(\omega_{d}t),$ (37)
$\dot{\sigma}_{j}^{-}={}$ $-\left[(\gamma_{j}+\gamma^{j}_{\rm loss})/2+i\omega_{j}^{q}\right]\sigma_{j}^{-}+ig_{j}(a+a)^{{\dagger}}\sigma_{j}^{z}$
$+\sqrt{\gamma_{j}}\sigma_{j}^{z}b_{\rm in}^{j}+\sqrt{\gamma^{j}_{\rm loss}}\sigma_{j}^{z}c_{\rm in}^{j}-i\sqrt{2\gamma_{\phi}^{j}}\chi(t)\sigma_{j}^{-},$ (38)
$\dot{\sigma}_{j}^{z}={}$ $-(\gamma_{j}+\gamma^{j}_{\rm loss})(1+\sigma_{j}^{z})-2ig_{j}(a^{{\dagger}}+a)(\sigma_{j}^{+}-\sigma_{j}^{-})$
$-2\sqrt{\gamma_{j}}\left[\sigma_{j}^{+}b_{\rm in}^{j}+{\rm h.c.}\right]-2\sqrt{\gamma^{j}_{\rm loss}}\left[\sigma_{j}^{+}c_{\rm in}^{j}+{\rm h.c.}\right].$ (39)
@@
$\delta\dot{a}={}$ $-\left[\kappa/2+i\omega_{c}\right]\delta a-i\sum_{j=1}^{N}g_{j}(\sigma_{j}^{-}+\sigma_{j}^{+})-\sqrt{\kappa}a_{\rm in}(t),$ (40)
@@
$\dot{\sigma}_{j}^{-}={}$ $-\left[(\gamma_{j}+\gamma^{j}_{\rm loss})/2+i\omega_{j}^{q}\right]\sigma_{j}^{-}+ig_{j}(\delta a+\delta a^{{\dagger}})\sigma_{j}^{z}$
$+2ig_{j}{\rm Re}\{\alpha\}\sigma_{j}^{z}+\sqrt{\gamma_{j}}\sigma_{j}^{z}b_{\rm in}^{j}+\sqrt{\gamma^{j}_{\rm loss}}\sigma_{j}^{z}c_{\rm in}^{j}$
$-i\sqrt{2\gamma_{\phi}^{j}}\Theta(t)\sigma_{j}^{-},$ (41)
$\dot{\sigma}_{j}^{z}={}$ $-(\gamma_{j}+\gamma_{\rm loss}^{j})(1+\sigma_{j}^{z})-2ig_{j}(\delta a^{{\dagger}}+\delta a)(\sigma_{j}^{+}-\sigma_{j}^{-})$
$-4ig_{j}{\rm Re}\{\alpha\}(\sigma_{j}^{+}-\sigma_{j}^{-})-2\sqrt{\gamma_{j}}\left[\sigma_{j}^{+}b_{\rm in}^{j}(t)+{\rm h.c.}\right]$
@@
Here, $\alpha(t)$ is determined via the classical differential equation (8).
@@
$H^{\prime}(t)={}$ $\omega_{c}\delta a^{{\dagger}}\delta a+\sum_{j=1}^{N}\frac{\omega_{j}^{q}}{2}\sigma_{j}^{z}+\sum_{j=1}^{N}g_{j}(\delta a+\delta a^{{\dagger}})(\sigma_{j}^{-}+\sigma_{j}^{+})$
$+\sum_{j=1}^{N}2g_{j}{\rm Re}\{\alpha(t)\}(\sigma_{j}^{+}+\sigma_{j}^{-}).$ (44)
Notice that the master equation (43) for the dynamics of $\delta a$ is the same as the master equation (6) for the total field $a$, except for the driving term in displaced Hamiltonian (44), which acts on the qubits rather than on the cavity as indicated in Eq. (10). This is the effective cavity-mediated driving on the qubits that allows for the synchronization of many SPSs in our scheme.
The displaced master equation (43) is very efficient to perform numerical simulations of the dynamics in the case of strong cavity driving $|\Omega(t)|\gg\kappa,|\Delta|$ because most of the cavity photons can be taken into account by the classical field $\alpha(t)$, and the fluctuation $\langle\delta a^{\dagger}\delta a\rangle$ is only weakly populated. This largely reduces the Hilbert space dimension needed to numerically simulate the dynamics of the driven cavity mode coupled to all qubits. In addition, if we consider the cavity in the weak coupling and bad cavity limit given by Eq. (11), the dynamics of the system can be further simplified by applying the rotating wave approximation (RWA) to the Hamiltonian in Eq. (44). On the one hand, since $g_{j}\ll\omega_{c}+\omega_{q}^{j}$ and the fluctuation is weakly populated, we can approximate the qubit-fluctuation coupling as
$g_{j}(\delta a+\delta a^{{\dagger}})(\sigma_{j}^{-}+\sigma_{j}^{+})\approx g_{j}(\delta a^{\dagger}\sigma_{-}+{\rm h.c.}).$ (45)
On the other hand, the cavity-mediated driving on the qubits can also be simplified by applying RWA as
$2g_{j}{\rm Re}\{\alpha(t)\}(\sigma_{j}^{+}+\sigma_{j}^{-})\approx g_{j}|\alpha_{\rm ss}|f(t)(\sigma_{j}^{+}e^{-i(\omega_{d}t-\phi)}+{\rm h.c.}).$ (46)
Here, $|\alpha_{\rm ss}|=\Omega_{0}/\sqrt{(\kappa/2)^{2}+\Delta^{2}}$ and $\phi={\rm Arctan}(2\Delta/\kappa)-\pi/2$ are the steady state amplitude and phase of the cavity displacement, respectively, and $f(t)$ is a step function profile. To demonstrate this last approximation (46), we first assume the square pulse is switched on $f(t)=1$ and the cavity is initially empty $\alpha(0)=0$. Then, the solution of the classical equation (8) for $\alpha(t)$ reads
$\alpha(t)={}$ $|\alpha_{\rm ss}|e^{-i(\omega_{d}t-\phi)}(1-e^{-(\kappa/2-i\Delta)t})$
$-\frac{\Omega_{0}}{2\omega_{d}}e^{i\omega_{d}t}(1-e^{-(\kappa/2+2i\omega_{d})t}),$ (47)
where we have also used the inequalities $|\Delta|,\kappa\ll\omega_{d}$ valid for our parameter conditions. The first term in (47) reaches the steady state value $\sim|\alpha_{\rm ss}|e^{-i(\omega_{d}t-\phi)}$ in a timescale $\sim 1/\kappa$ and the second term causes the fast but low amplitude oscillation observed in Fig. 2(a). Even for a large driving strength $\Omega_{0}\gtrsim\omega_{d}$, the second term can be neglected with respect to the first when $|\alpha_{\rm ss}|\gg 1$ as it is the case for our parameters. Since in bad cavity limit (11) the timescale for the qubit dynamics is much longer than $\sim 1/\kappa$, we can use the steady state value $\alpha(t)\approx|\alpha_{\rm ss}|e^{-i(\omega_{d}t-\phi)}$ to approximate the cavity displacement when the drive is switched on $f(t)=1$ as it would be instantaneous for the qubits. Similarly, after the cavity reaches the steady state and it is switched off ($f(t)=0$), the dynamics in Eq. (8) predicts an exponential decay of $\alpha(t)$ with rate $\sim\kappa$. In the bad cavity limit we can also approximate this emptying of the cavity as instantaneous for the qubit and thus $\alpha(t)\approx 0$ for $f(t)=0$.
In summary, under the above approximations we have $\alpha(t)\approx|\alpha_{\rm ss}|e^{-i(\omega_{d}t-\phi)}f(t)$ so that when replacing it in the last term of Eq. (44), and applying RWA provided $g_{j}|\alpha_{\rm ss}|\ll\omega_{d}$, we obtain Eq. (46). Replacing Eqs. (45)-(46) in (44), and going to a rotating frame with respect to the cavity drive frequency $\omega_{d}$, the dynamics of the system in the displaced picture is finally given by the master equation (43) with the RWA Hamiltonian:
$H^{\prime}(t)\approx{}$ $-\Delta\delta a^{{\dagger}}\delta a-\frac{1}{2}\sum_{j=1}^{N}\delta_{j}\sigma_{j}^{z}+\sum_{j=1}^{N}g_{j}(\delta a^{{\dagger}}\sigma_{j}^{-}+\sigma_{j}^{+}\delta a)$
$+\sum_{j=1}^{N}\Omega_{\rm cm}^{j}(\sigma_{j}^{+}+\sigma_{j}^{-}).$ (48)
Here, we have defined $\Omega_{\rm cm}^{j}=g_{j}|\alpha_{\rm ss}|$ as in Eq. (14) and absorbed the constant phase $\phi$ in the definition of the qubit and fluctuation operators. The qubit detunings read $\delta_{j}=\omega_{d}-\omega_{q}^{j}$ as in the main text.
@@ A.2 Adiabatic elimination of cavity fluctuations
As discussed in the previous subsection, in the bad cavity limit $g_{j},|\delta_{j}|\ll\kappa$, the cavity reaches very quickly a coherent steady state, with its quantum fluctuations $\delta a$ close to the vacuum state $|0\rangle$. In this situation it is convenient to adiabatically eliminate the cavity fluctuation and obtain an effective dynamics for the degrees of freedom of the qubits only. To do so, we formally integrate Eq. (40) and apply Markov approximation provided the qubits’ dynamics evolves slowly on the time-scale $\sim 1/\kappa$. As a result, we get
$\delta a(t)\approx{}$ $-i\sum_{j}\frac{g_{j}}{(\kappa/2-i\Delta)}\sigma_{j}^{-}(t)-\frac{\sqrt{\kappa}}{(\kappa/2-i\Delta)}a_{\rm in}(t)$
${\color[rgb]{0,0,0}-i\sum_{j}\frac{g_{j}}{\kappa/2+i(2\omega_{d}-\Delta)}\sigma_{j}^{+}(t).}$ (49)
Replacing the above expressions in the original quantum Langevin equations (38)-(39), applying RWA, and going to a rotating frame with frequency $\omega_{d}$, we obtain the effective dynamics of the qubits, given by the Langevin equations
@@
$-2\sqrt{\gamma^{j}_{\rm loss}}\left[\sigma_{j}^{+}c_{\rm in}^{j}+{\rm h.c.}\right].$ (51)
Here, the effective cavity-mediated decay $\gamma_{\rm cm}^{j}$, detuning $\delta_{\rm cm}^{j}$, hopping $J_{\rm cm}^{jl}$, and driving strength $\Omega_{\rm cm}^{j}$ are defined in Eqs. (14)-(17).
Finally, we can find a master equation which leads to the same effective Langenvin equations in Eqs. (50)-(51). This effective master equation, which incorporates the adiabatic elimination of the cavity mode, is given in Eqs. (12)-(13).
@@ Appendix B Photon counting and calculation of photon generation probabilities
Photon detection and counting lies at the heart of Quantum Optics and thus plenty of methods have been developed over time Gardiner and Zoller (2004); Plenio and Knight (1998). In this Appendix we describe two methods for quantifying the efficiencies and photon statistics from the output of many SPSs. In Sec. B.1 we introduce an original photon counting method that we developed based on extending the master equation formalism. Then, in Sec. B.2, we discuss a more standard photon counting method using the quantum trajectory (QT) approach. This is useful when the Hilbert space of the system becomes very large, but at the expense of losing precision in the computation of the averages.
@@ B.2 Photon counting in quantum trajectories formalism
The most natural way to implement an ideal photon counting is within the formalism of quantum trajectories (QT) and continuous measurements Plenio and Knight (1998); Daley (2014); Dalibard et al. (1992); Mølmer et al. (1993); Dum et al. (1992). Here, the physics of quantum jumps is explicitly simulated during the open system evolution and therefore it is very natural to count them and thereby infer the photon statistics.